OpenStudy (anonymous):

I can't seem to understand what this question is asking for. Help? Use differentials to approximate the quantity. (Give your answer correct to 4 decimal places.) √10 (there's supposed to be a line over the 10 breaking the radical, which I don't know what that means.)

OpenStudy (amistre64):

can you draw it with the draw button?

OpenStudy (amistre64):

the OS is very jittery today :/

OpenStudy (inkyvoyd):

We are using the newton raphson iteration method.

OpenStudy (anonymous):



OpenStudy (inkyvoyd):

x1=x0-(f(x0)/f'(x0))

OpenStudy (inkyvoyd):

If this is what they want, keep looking. If it isn't, then ignore me.

OpenStudy (experimentx):

(10)^(1/2) = (1+9)^(1/2) = 3(1+1/9)^(1/2) 3{1+1/2*1/9-1/2*1/2*1/2!*(1/9)^2 - 3/2*1/2*1/2*1/3!*1/9^3}

OpenStudy (anonymous):

this is part of the Differentials section of my Business Calc class..

OpenStudy (inkyvoyd):

x1=x0-(x0^2)/(2x0)

OpenStudy (inkyvoyd):